# MicroAdam: Accurate Adaptive Optimization with Low Space Overhead and Provable Convergence

### v1

@@ 1 Introduction
In this paper, we address this gap by introducing MicroAdam, an adaptive optimizer which guarantees low memory usage but also ensures provable convergence. We start from the idea that we can allow the (lossy) sparse projection of gradient information before it enters the optimizer states; crucially, different from prior work, we ensure convergence by correcting for the inherent error due to compression by employing a novel variant of error correction, a mechanism introduced for distributed optimization (Seide et al., 2014). However, simply using error feedback would not lead to memory savings, since the size of the error correction buffer is comparable to the that of the original optimizer state. Instead, our main algorithmic innovation is in showing that the error feedback can itself be compressed in the context of adaptive optimization. This renders the memory overhead of error feedback negligible, while preserving convergence guarantees.
Specifically, on the theoretical side, we provide a new analysis showing that, under reasonable assumptions on the loss function being optimized and on the degree of compression, MicroAdam provably guarantees convergence, at asymptotically the same rate as AMSGrad (Zhou et al., 2024a), i.e. a version of Adam with general convergence guarantees. The key finding is that our approach allows for the overhead of compression to be shifted to the higher-order terms, where it should not impact practical convergence in common cases. This claim holds both for general smooth non-convex functions, and for non-convex functions under the Polyak-Lojasiewicz (PL) assumption, highlighting a trade-off between the degree of compression of the gradients, and that of the error feedback.
We complement our algorithmic and analytic results with an efficient GPU implementation of MicroAdam, which we validate for fine-tuning language models from the BERT (Devlin et al., 2018), OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023) families, with hundreds of millions to billions of parameters. We show that, in practice, gradients can be projected to very high sparsity (99%), while the error correction can be stored at 4 bits per component, without loss of convergence. Concretely, our method can significantly improve upon the memory footprint of the extremely popular 8bit Adam (Dettmers et al., 2021) when fine-tuning models such as LLaMA2-7B (Touvron et al., 2023), at similar or better accuracy. At the same time, MicroAdam provides better accuracy relative to high-compression heuristics such as GaLore (Zhao et al., 2024).
@@ 2 Related Work
Due to space constraints, we mainly focus on related work reeducing the cost of optimizer states. Dettmers et al. (2021) considers this problem, specifically by performing fine-grained quantization of the optimizer states. Their work does not alter the Adam algorithm; instead, it deals with the challenge of accurately compressing the dynamically-changing meta-data sequence. As the name suggests, the space savings correspond to roughly halving the memory required by the optimizer states, relative to FP16. In the same vein, AdaFactor (Shazeer and Stern, 2018) and CAME (Luo et al., 2023) reduce memory cost by factorizing the second-order statistics, while the recent GaLore (Zhao et al., 2024) factorizes the gradients themselves before they enter the optimizer state (but does not use error correction). Importantly, these methods are heuristics: they do not provide theoretical guarantees under standard assumptions, [^1] GaLore (Zhao et al., 2024) does state convergence guarantees for a variant of the algorithm with fixed projections, but this is under a strong “stable rank” assumption, which may not hold in practice. and in practice require careful tuning to preserve convergence (Luo et al., 2023). By contrast, our method is theoretically justified, and provides good practical convergence. Earlier work by Anil et al. (2019) provides convergence guarantees for a compressed variant of Adagrad (Duchi et al., 2010) called SM3, improving upon earlier work by Spring et al. (2019). However, it is not clear how to extend their approach to the popular Adam optimizer, and heuristic methods appear to provide superior performance (Luo et al., 2023).
@@ 3 The MicroAdam Algorithm
Algorithm 1 MicroAdam: Implementation View
1:Input: $\beta_{1},\beta_{2},\epsilon,\mathcal{G},T,d$
2:$m_{0},v_{0},e_{1},\delta_{1},\Delta_{1}\leftarrow 0_{d},0_{d},0_{d}^{4b},0_{d},0_{d}$
3:for $t=\{1,2,...,T\}$ do
4:     $a_{t}\leftarrow\widetilde{\nabla}_{\theta}f(\theta_{t})+{\color[rgb]{0,0,1}Q^{-1}(e_{t},\delta_{t},\Delta_{t}})$
5:     $\mathcal{I}_{t},\mathcal{V}_{t}\leftarrow T_{k}(|a_{t}|)$
6:     $a_{t}[\mathcal{I}_{t}]\leftarrow 0$
7:     $\delta_{t+1},\Delta_{t+1}\leftarrow min(a_{t}),max(a_{t})$
8:     $e_{t+1}\leftarrow Q(a_{t},\delta_{t+1},\Delta_{t+1})$
9:     $\mathcal{G}_{i,:}\leftarrow(\mathcal{I}_{t},\mathcal{V}_{t})$
10:     $\hat{m_{t}}\leftarrow\textsc{AdamStats}(\beta_{1},\mathcal{G})$
11:     $\hat{v_{t}}\leftarrow\textsc{AdamStats}(\beta_{2},\mathcal{G}^{2})$
12:     $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\frac{\hat{m_{t}}}{\epsilon+\sqrt{\hat{v_{t}}}}$
13:     $i\leftarrow(i+1)\%m$
14:end for
@@
In Algorithm 3, in line 4 we dynamically re-compute the decay factor $\beta^{r}$ with exponent $r$ determined based on the current optimization step $t$, $i^{th}$ row of the buffer $\mathcal{G}$ and the window size $m$. The most recent row will have $r=0$, while the latest one will have $r=m-1$. In the end, we will add the values $\beta^{r}\cdot\mathcal{V}_{i}$ to the buffer $z$ at the corresponding indices $\mathcal{I}_{i}$. We discuss the efficient CUDA implementation in the Appendix.
Algorithm 2 Symmetric Uniform Quantization and Dequantization
@@
Algorithm 3 Maintaining Adam Statistics
1:procedure AdamStats($\beta,\mathcal{G},t,m,d$)
2:     $z\leftarrow 0_{d}$
3:     for $i\in\{1,2,...,\min(t,m)\}$ do
4:         $r\leftarrow(t-i-1)\%m$
5:         $z[\mathcal{I}_{i}]\leftarrow z[\mathcal{I}_{i}]+\beta^{r}\cdot\mathcal{V}_{i}$
6:     end for
7:     return $z/(1-\beta^{t})$
8:end procedure
In detail, at each optimization step $t$ we add up the current gradient $g_{t}=\nabla_{\theta}f(\theta_{t})$ and the dequantized error (see procedure $Q^{-1}$ in Algorithm 2) from the previous iterations (feeding back the error) into the accumulator variable $a_{t}$. The accumulator is compressed using the Top-K operator before being passed to Adam. The pair $(\mathcal{I}_{t},\mathcal{V}_{t})$ represents the indices of the largest components by absolute value and the corresponding values (this step is equivalent to selecting the outliers from $a_{t}$). In step 6, we remove the outliers from the accumulator $a_{t}$ to finally obtain the error at the current step $t$, which is still stored in $a_{t}$. From an algorithmic perspective, this step is equivalent to writing $e_{t}\leftarrow a_{t}-T_{k}(a_{t})$.
Next, we compress the error feedback. In our practical implementation, we will use compression to 4-bit integers, more specifically using symmetric uniform quantization in large blocks of fixed size $B_{q}$ (step 8). More precisely, this computes the minimum and maximum values $\delta_{t}$ and $\Delta_{t}$ (step 7) to quantize the accumulator $a_{t}$, as detailed in the procedure $Q$.
Finally, we update the $i^{th}$ row in the buffer $\mathcal{G}$ by replacing the oldest pair of indices and values with the one newly computed at step $t$. The statistics $m_{t}$ and $v_{t}$ are both estimated using the sliding window $\mathcal{G}$, where $\mathcal{G}^{2}=(\mathcal{I},\mathcal{V}^{2})$ (via element-wise squaring). We would like to point out that the resulting update $u_{t}=m_{t}/(\epsilon+\sqrt{v_{t}})$ may be highly sparse when the window size $m$ is small. For illustration, if we use density $k=d/100$ ($1\%$ density, e.g. $99\%$ sparsity) with $m=10$ and suppose that all rows in the indices matrix $\mathcal{I}$ are disjoint, then the overall density in the update $u_{t}$ will be at least $90\%$. (Note that the sparsity increases if rows in $\mathcal{I}$ have common values.) However, over large windows, the update sparsity should decrease.
Algorithm Intuition.
To gain more intuition, we illustrate the impact of compressing gradients using TopK compression before they are incorporated into the optimizer state for Adam, both with and without error feedback (EF). We use the classic 2D Rosenbrock function $f(x,y)=(1-x)^{2}+100(y-x^{2})^{2}$ to illustrate this effect. In the Appendix, we provide similar results for another ill-conditioned problem. The optimization trajectories are discussed in Figure 1. The results clearly show that EF is essential for fast convergence. Besides, TopK with EF, which is a surrogate of MicroAdam, allows for competitive convergence relative to the uncompressed baseline. In Appendix E we discuss implications of Error Feedback applied to GaLore.
Figure 1: Optimization trajectories of Adam, TopK-Adam and TopK-Adam with EF applied on the Rosenbrock function starting from $(x_{0},y_{0})=(-\frac{1}{2},1)$. Notice the extremely “jagged” profile of TopK-Adam without EF, and the recovered convergence when EF is added.
@@ 3.1 Efficient Implementation
First, we do not use an additional accumulator tensor $a_{t}$; instead, we use a CUDA kernel to dequantize the error buffer, and store the result in the grad attribute of the model parameters. This allows us to accumulate the error feedback into the gradients, without any additional memory.
@@
Since we run on LLMs with billions of parameters, naive storage of the sparse indices would require using an int64 type for the indices matrix $\mathcal{I}$, assuming that the Top-K operator is applied globally to all the parameters in $a_{t}$. To avoid this cost, we apply Top-K in blocks of fixed size $B_{d}<2^{16}-1=65535$ and store block-relative indices in int16 format. Applying Top-K per row to $a_{t}$ reshaped to 2D is not only faster, but provides the block-relative indices directly.
Computing Top-K in blocks also allows us to allocate and efficiently use CUDA shared memory blocks to compute the statistics $m_{t}$ and $v_{t}$ for Adam, as described in the AdamStats procedure below. Note that the block-relative indices returned by Top-K will be directly used as indices in the shared memory array of CUDA thread blocks to retrieve values from $\mathcal{I}$ and $\mathcal{V}$.
@@
3.2 Memory footprint analysis and comparison with other methods
We now compare the theoretical memory footprint of MicroAdam AdamW (Loshchilov and Hutter, 2019) and AdamW-8 bits (Dettmers et al., 2021), and GaLore (Zhao et al., 2024), focusing on memory usage of the optimizer states $m_{t}$ and $v_{t}$, each of size $d$, expressed in bytes (B). For concreteness, we report the practical memory usage for the optimizer state for a Llama-2 7B model for each optimizer.
@@ Discussion.
Assume that our goal is to obtain a lower memory footprint compared to AdamW-8 bit. Recall that we use sparsification to compress gradients; then, we will compute the gradient density $k=d/100$ for MicroAdam in order to be competitive with AdamW-8 bit.
For this, we have to solve the equation $M_{\mu A}(m)=M_{AW8}$ for $m$, resulting in $m_{\max}=37.5$. Specifically, if we use a gradient history of $m<m_{\max}$ gradients in $\mathcal{G}$, MicroAdam will have theoretical memory savings. We will see that, in practice, this history size is more than enough for good practical results. As entries in the window past this range are dampened extremely significantly, their influence is negligible. In Appendix C, we provide Python code to compute the theoretical memory usage for the three optimizers MicroAdam for Llama-2 7B.

### v2

@@ 1 Introduction
In this paper, we address this gap by introducing MicroAdam, an adaptive optimizer which guarantees low memory usage but also ensures provable convergence. We develop our approach to improve the performance of finetuning LLMs and mainly focus on the research question “are all gradient entries important for optimization?” To answer this question, we start from the idea that we can allow the (lossy) sparse projection of gradient information before it enters the optimizer states; crucially, different from prior work, we ensure convergence by correcting for the inherent error due to compression by employing a novel variant of error correction, a mechanism introduced for distributed optimization (Seide et al., 2014). However, simply using error feedback would not lead to memory savings, since the size of the error correction buffer is comparable to the that of the original optimizer state. Instead, our main algorithmic innovation is in showing that the error feedback can itself be compressed in the context of adaptive optimization. This renders the memory overhead of error feedback negligible, while preserving convergence guarantees.
Specifically, on the theoretical side, we provide a new analysis showing that, under reasonable assumptions on the loss function being optimized and on the degree of compression, MicroAdam provably guarantees convergence, at asymptotically the same rate as AMSGrad (Zhou et al., 2024a), i.e. a version of Adam with general convergence guarantees, that fixes a fundamental technical issue in the Adam optimizer’s proof  (Reddi et al., 2019). The key finding is that our approach allows for the overhead of compression to be shifted to the higher-order terms, where it should not impact practical convergence in common cases. This claim holds both for general smooth non-convex functions, and for non-convex functions under the Polyak-Lojasiewicz (PL) assumption, highlighting a trade-off between the degree of compression of the gradients, and that of the error feedback.
We complement our algorithmic and analytic results with an efficient GPU implementation of MicroAdam, which we validate for fine-tuning language models from the BERT (Devlin et al., 2018), OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023) families, with hundreds of millions to billions of parameters. We show that, in practice, gradients can be projected to very high sparsity (99%), while the error correction can be stored at 4 bits per component, without loss of convergence. Concretely, our method can significantly improve upon the memory footprint of the extremely popular 8bit Adam (Dettmers et al., 2021) when fine-tuning models such as LLaMA2-7B/13B (Touvron et al., 2023), at similar or better accuracy. At the same time, MicroAdam provides better accuracy relative to high-compression heuristics such as GaLore (Zhao et al., 2024).
@@ 2 Related Work
We mainly focus on related work reducing the cost of optimizer states. Dettmers et al. (2021) considers this problem, specifically by performing fine-grained quantization of the optimizer states. Their work does not alter the Adam algorithm; instead, it deals with the challenge of accurately compressing the dynamically-changing meta-data sequence. As the name suggests, the space savings correspond to roughly halving the memory required by the optimizer states, relative to FP16. In the same vein, AdaFactor (Shazeer and Stern, 2018) and CAME (Luo et al., 2023) reduce memory cost by factorizing the second-order statistics, while the recent GaLore (Zhao et al., 2024) factorizes the gradients themselves before they enter the optimizer state (but does not use error correction). Importantly, these methods are heuristics: they do not provide theoretical guarantees under standard assumptions, [^1] GaLore (Zhao et al., 2024) does state convergence guarantees for a variant of the algorithm with fixed projections, but this is under a strong “stable rank” assumption, which may not hold in practice. and in practice require careful tuning to preserve convergence (Luo et al., 2023). By contrast, our method is theoretically justified, and provides good practical convergence. Earlier work by Anil et al. (2019) provides convergence guarantees for a compressed variant of Adagrad (Duchi et al., 2010) called SM3, improving upon earlier work by Spring et al. (2019). However, it is not clear how to extend their approach to the popular Adam optimizer, and heuristic methods appear to provide superior performance (Luo et al., 2023).
@@ 3 The MicroAdam Algorithm
In detail, at step $t=1$, the error feedback $e_{1}$ is completely zero, as initialized in line 2, and thus, at line 5 the accumulator $a_{1}$ will only contain the stochastic gradient $g_{1}$. At line 6, we perform the Top-K compression and only keep the top-$1\%$ of values $\mathcal{V}_{1}$ and their corresponding indices $\mathcal{I}_{1}$. The compression is equivalent to choosing the top-$1\%$ values in the left and right tails (outliers) due to the absolute value we apply on top of accumulator $a$. At line 7, we remove the outliers from the accumulator because they will be transferred to the buffer matrix $\mathcal{G}$. This step is equivalent to $e\leftarrow a-T_{k}(a)$ found in theoretical results. After line 7, what is left in $a$ is called the error feedback (e.g. the weights which were not chosen by Top-k). At line 8, we compute the statistics $\delta$ and $\Delta$ needed for quantization, and at line 9, we effectively quantize the accumulator (e.g. error feedback after line 7). At line 10 we update the buffer $\mathcal{G}$, in lines 11, 12 and 13 we compute the statistics $\hat{m},\hat{v}$ (computed by squaring the entries of $\mathcal{G}$ element-wise) and update the model parameters.
For steps $t\geq 2$, the only change compared to $t=1$ is that error feedback $e$ is not zero anymore. Since the error is compressed, we need to decompress it and add it to the gradient. This process happens at line 5 and it is the point where we feed back the error: the accumulator will store the gradient whose direction is corrected by the error (e.g. the cumulative history of weights not chosen by Top-k at the previous steps).
Properties and Limitations.
We would like to point out that the resulting update $u_{t}=m_{t}/(\epsilon+\sqrt{v_{t}})$ will always be highly sparse when the window size $m$ is small. For illustration, if we use density $k=d/100$ (e.g. $1\%$ density equivalent to $99\%$ sparsity) with $m=10$ and suppose that all rows in the indices matrix $\mathcal{I}$ are disjoint, then the overall density in the update $u_{t}$ will be $90\%$. The sparsity of $u_{t}$ increases if rows in $\mathcal{I}$ have common values. MicroAdam yields good results for LLM finetuning and pre-training computer vision models, as the experimental section shows. However, we noticed the update $u_{t}$ of MicroAdam is too sparse to be able to provide good enough updates for LLM pre-training. We believe this happens because the attention layers must receive dense updates to be able to learn the correlations between words.
@@
In AdamStats procedure in Algorithm 2 we implement the unrolled recursion of momentum $z_{t}\leftarrow\beta z_{t-1}+(1-\beta)g_{t}$ for the last $m$ sparse gradients as $z_{t}\leftarrow(1-\beta)\sum_{i=t-m}^{t}\beta^{t-i}g_{i}$ and we also perform the bias correction in the end. Because we compute $\hat{m}_{t}$ and $\hat{v}_{t}$ using the last $m$ sparse gradients in the window, in line 4 we dynamically determine the exponent $r$ for the decay factor $\beta^{r}$ based on the current optimization step $t$, $i^{th}$ row of the circular buffer $\mathcal{G}$ and the window size $m$. The last gradient added to $\mathcal{G}$ will have $r=0$, while the oldest gradient in $\mathcal{G}$ will have $r=m-1$. In the end, we will add the values $\beta^{r}\mathcal{V}_{i}$ to the buffer $z$ at the corresponding indices $\mathcal{I}_{i}$, which is a fast operation because we only manipulate $1\%$ of values at a time. We discuss the efficient CUDA implementation in the Appendix.
Algorithm Intuition.
To gain intuition, we illustrate the impact of compressing gradients via TopK before they are incorporated into the optimizer state for Adam, both with and without error feedback (EF). Figure 1 shows how EF fixes AdamW with Top-K compression. The plot on the left shows the optimization trajectory of the original Adam optimizer. The center plot illustrates the convergence of Top-K Adam when we only choose the largest coordinate from the accumulator (equivalent to $50\%$ sparsity since the problem is 2D). In the end, on the right side we show that adding EF to Top-K Adam recovers the same optimization trajectory as the original Adam optimizer. Extrapolating to higher dimensional problems, our MicroAdam approach helps recover the trajectory of the original Adam optimizer, while using less memory. The results clearly show that EF is essential for fast convergence. Besides, TopK with EF, which is a surrogate of MicroAdam, allows for competitive convergence relative to the uncompressed baseline. In Appendix F, we discuss the implications of Error Feedback applied to GaLore.
Algorithm 1 Pseudocode for MicroAdam with quantized EF
1:Input: $\beta_{1},\beta_{2},\epsilon,\mathcal{G},T,d,k$
2:$m_{0},v_{0}\leftarrow 0_{d},0_{d}$
3:$\delta_{1},\Delta_{1}\leftarrow 0,0$
4:$e_{1}\leftarrow 0_{d}^{4b}$
5:for $t=\{1,2,...,T\}$ do
6:     $g_{t}\leftarrow\widetilde{\nabla}_{\theta}f(\theta_{t})$
7:     $a_{t}\leftarrow g_{t}+{\color[rgb]{0,0,1}Q^{-1}(e_{t},\delta_{t},\Delta_{t}})$
8:     $\mathcal{I}_{t},\mathcal{V}_{t}\leftarrow T_{k}(|a_{t}|)$
9:     $a_{t}[\mathcal{I}_{t}]\leftarrow 0$
10:     $\delta_{t+1},\Delta_{t+1}\leftarrow min(a_{t}),max(a_{t})$
11:     $e_{t+1}\leftarrow Q(a_{t},\delta_{t+1},\Delta_{t+1})$
12:     $\mathcal{G}_{i,:}\leftarrow(\mathcal{I}_{t},\mathcal{V}_{t})$
13:     ${\color[rgb]{0,0,0}\hat{m}_{t}}\leftarrow\textsc{AdamStats}(\beta_{1},\mathcal{G})$
14:     ${\color[rgb]{0,0,0}\hat{v}_{t}}\leftarrow\textsc{AdamStats}(\beta_{2},\mathcal{G}^{2})$
15:     $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\frac{{\color[rgb]{0,0,0}\hat{m}_{t}}}{\epsilon+\sqrt{{\color[rgb]{0,0,0}\hat{v}_{t}}}}$
16:     $i\leftarrow(i+1)\%m$
17:end for
Algorithm 2 Adam Statistics, Quantization and Inverse Quantization
1:procedure AdamStats($\beta,\mathcal{G},t,m,d$)
2:     $z\leftarrow 0_{d}$
3:     for $i\in\{1,2,...,\min(t,m)\}$ do
4:         $r\leftarrow(t-i-1)\%m$
5:         $z[\mathcal{I}_{i}]\leftarrow z[\mathcal{I}_{i}]+\beta^{r}\mathcal{V}_{i}$
6:     end for
7:     return $\frac{{\color[rgb]{0,0,0}(1-\beta)z}}{1-\beta^{t}}$
8:end procedure
@@
Figure 1: Optimization trajectories of Adam, TopK-Adam and TopK-Adam with EF applied on the Rosenbrock function $f(x,y)=(1-x)^{2}+100(y-x^{2})^{2}$ starting from $(x_{0},y_{0})=(-\frac{1}{2},1)$. Notice the extremely “jagged” profile of TopK-Adam without EF, and the recovered convergence when EF is added.
@@ 3.1 Efficient Implementation
First, we do not use an additional accumulator tensor $a_{t}$; instead, we use a CUDA kernel to dequantize the error buffer, and store the result in the grad attribute of the model parameters. This allows us to accumulate the error feedback into the gradients, without allocating a full or half precision $d$-dimensional array. Each component of the EF has 4 bits and the entire EF is stored in an array of size $d/2$ of uint8 values.
@@
Since we run on LLMs with billions of parameters, naive storage of the sparse indices would require using an int64 type for the indices matrix $\mathcal{I}$, assuming that the Top-K operator is applied globally to all the parameters in $a_{t}$. To avoid this cost, we apply Top-K in blocks of fixed size $B_{d}<2^{15}-1=32767$ and store block-relative indices in int16 format (during the development of MicroAdam, PyTorch did not have support for uint16). Applying Top-K per row to $a_{t}$ reshaped to 2D is not only faster, but provides the block-relative indices directly.
Computing Top-K in blocks also allows us to allocate and efficiently use CUDA shared memory blocks to dynamically compute the statistics $\hat{m}_{t}$ and $\hat{v}_{t}$ for Adam, as described in the AdamStats procedure in Algorithm 2. We allocate the maximum possible shared memory for each thread block and store $\hat{m}_{t}$ (first half) and $\hat{v}_{t}$ (second half) at consecutive locations in the shared memory. Once these statistics are computed, it is easy to update the model parameters. Note that the block-relative indices returned by Top-K will be directly used as indices in the shared memory array of CUDA thread blocks to retrieve values from $\mathcal{I}$ and $\mathcal{V}$.
@@
3.2 Memory footprint analysis for the optimizer states and comparison with other methods
We now compare the theoretical memory footprint of MicroAdam with AdamW (Loshchilov and Hutter, 2019), AdamW-8 bits (Dettmers et al., 2021), and GaLore (Zhao et al., 2024), focusing on memory usage of the optimizer states $m_{t}$ and $v_{t}$, each of size $d$, expressed in bytes (B). For concreteness, we report the practical memory usage for the optimizer state for a Llama-2 7B model for each optimizer.
@@ Discussion.
Assume our goal is to obtain a lower memory footprint compared to AdamW-8 bit. We fix the gradient density to $k=d/100$ and we have to determine the number of gradients (window size) $m$ for MicroAdam in order to be competitive with AdamW-8 bit.
For this, we have to solve the equation $M_{\mu A}(m)=M_{AW8}$ for $m$, resulting in $m_{\max}=37.5$. Specifically, if we use a gradient history of $m<m_{\max}$ gradients in $\mathcal{G}$, MicroAdam will have theoretical memory savings. We will see that, in practice, this history size $m=10$ is more than enough for good practical results. As entries in the window past this range are dampened extremely significantly, their influence is negligible. In Appendix D, we provide Python code to compute the theoretical memory usage for the three optimizers for Llama-2 7B.